Given the function f, discuss its relative maximum and minimum

points, the intervals where it is increasing and decreasing, the intervals of concavity, and the points of inflection. Construct a sketch of the graph of the function.


1. 𝑓(π‘₯)= 2π‘₯βˆ’4 / xΒ²


2. 𝑓 π‘₯ =10π‘₯ / 1+3π‘₯Β²


3. 𝑓 π‘₯ = π‘₯Β³ βˆ’ 3/2 π‘₯Β²


4. 𝑓 π‘₯ = π‘₯ βˆ’ π‘₯Β³ / 3


1
Expert's answer
2021-06-21T06:08:36-0400

1. "f(x)=2x-4\/x^2"


"Df: (-\infin, 0)\cup(0, \infin)"



"f'(x)=(2x-\dfrac{4}{x^2})'=2+\dfrac{8}{x^3}"

Critical number(s):


"f'(x)=0=>2+\dfrac{8}{x^3}=0=>x=-\sqrt[3]{4}"


"f(-\sqrt[3]{4})=2(-\sqrt[3]{4})-4/(-\sqrt[3]{4})^2=-3\sqrt[3]{4}"

If "x<-\sqrt[3]{4}, f'(x)>0, f(x)" increases.


If "-\sqrt[3]{4}<x<0, f'(x)<0, f(x)" decreases.


If "x>0, f'(x)>0, f(x)" increases.


The function "f(x)" increases on "(-\infin, -\sqrt[3]{4})\cup(0, \infin)."

The function "f(x)" decreases on "(-\sqrt[3]{4}, 0)."


The function "f(x)" has a local maximum at "x=-\sqrt[3]{4}."

The function "f(x)" has no local minimum.


"f''(x)=(2+\dfrac{8}{x^3})'=-\dfrac{32}{x^4}<0 x\in Df"

The function "f(x)" is concave down on "(-\infin, 0)\cup(0, \infin)."

The function "f(x)" is not concave up anywhere.

The function "f(x)" has no point of inflection.




2. "f(x)=\dfrac{10x}{1+3x^2}"


"Df: (-\infin, \infin)"



"f'(x)=(\dfrac{10x}{1+3x^2})'=-\dfrac{10(3x^2-1)}{(1+3x^2)^2}"

Critical number(s):


"f'(x)=0=>-\dfrac{10(3x^2-1)}{(1+3x^2)^2}=0=>x=\pm\dfrac{\sqrt{3}}{3}"


"f(-\dfrac{\sqrt{3}}{3})=\dfrac{10(-\dfrac{\sqrt{3}}{3})}{1+3(-\dfrac{\sqrt{3}}{3})^2}=-\dfrac{5\sqrt{3}}{3}"

"f(\dfrac{\sqrt{3}}{3})=\dfrac{10(\dfrac{\sqrt{3}}{3})}{1+3(\dfrac{\sqrt{3}}{3})^2}=\dfrac{5\sqrt{3}}{3}"


If "x<-\dfrac{\sqrt{3}}{3}, f'(x)<0, f(x)" decreases.


If "-\dfrac{\sqrt{3}}{3}<x<\dfrac{\sqrt{3}}{3}, f'(x)>0, f(x)" increases.


If "x>\dfrac{\sqrt{3}}{3}, f'(x)<0, f(x)" decreases.


The function "f(x)" increases on "(-\dfrac{\sqrt{3}}{3}, \dfrac{\sqrt{3}}{3})."

The function "f(x)" decreases on "(-\infin, -\dfrac{\sqrt{3}}{3})\cup(\dfrac{\sqrt{3}}{3}, \infin)."


The function "f(x)" has a local maximum at "x=\dfrac{\sqrt{3}}{3}."

The function "f(x)" has a local minimum at "x=-\dfrac{\sqrt{3}}{3}."



"f''(x)=(-\dfrac{10(3x^2-1)}{(1+3x^2)^2})'=\dfrac{180x(x^2-1)}{(1+3x^2)^3}"




"f''(x)=0=>\dfrac{180x(x^2-1)}{(1+3x^2)^3}=0"

"x_1=-1, x_2=0, x_3=1"

"f(-1)=\dfrac{10(-1)}{1+3(-1)^2}=-\dfrac{5}{2}"

"f(0)=\dfrac{10(0)}{1+3(0)^2}=0"

"f(1)=\dfrac{10(1)}{1+3(1)^2}=\dfrac{5}{2}"

The function "f(x)" is concave up on "(-1, 0)\cup(1, \infin)."

The function "f(x)" is concave down on "(-\infin, -1)\cup(0, 1)."


The points of inflection are "(-1, -2.5), (0,0), (1,2.5)."





3. "f(x)=x^3-\dfrac{3}{2x^2}"


"Df: (-\infin, 0)\cup(0, \infin)"


"f'(x)=(x^3-\dfrac{3}{2x^2})'=3x^2+\dfrac{3}{x^3}"

Critical number(s):


"f'(x)=0=>3x^2+\dfrac{3}{x^3}=0=>x=-1"


"f(-1)=(-1)^3-\dfrac{3}{2(-1)^2}=-\dfrac{5}{2}"

If "x<-1, f'(x)>0, f(x)" increases.


If "-1<x<0, f'(x)<0, f(x)" decreases.


If "x>0, f'(x)>0, f(x)" increases.


The function "f(x)" increases on "(-\infin, -1)\cup(0, \infin)."

The function "f(x)" decreases on "(-1, 0)."


The function "f(x)" has a local maximum at "x=-1."

The function "f(x)" has no local minimum.


"f''(x)=(3x^2+\dfrac{3}{x^3})'=6x-\dfrac{9}{x^4}""f''(x)=0=>6x-\dfrac{9}{x^4}=0"

"x=\sqrt[5]{1.5}"

"f(\sqrt[5]{1.5})=(\sqrt[5]{1.5})^3-\dfrac{3}{2(\sqrt[5]{1.5})^2}=0"

The function "f(x)" is concave up on "(\sqrt[5]{1.5}, \infin)."

The function "f(x)" is concave down on "(-\infin, 0)\cup(0, \sqrt[5]{1.5})."


The point of inflection is "(\sqrt[5]{1.5},0)."



4. "f(x)=x-\dfrac{x^3}{3}"


"Df: (-\infin, \infin)"



"f'(x)=(x-\dfrac{x^3}{3})'=1-x^2"

Critical number(s):


"f'(x)=0=>1-x^2=0=>x=\pm1"


"f(-1)=-1-\dfrac{(-1)^3}{3}=-\dfrac{2}{3}"

"f(1)=1-\dfrac{(1)^3}{3}=\dfrac{2}{3}"


If "x<-1, f'(x)<0, f(x)" decreases.


If "-1<x<1, f'(x)>0, f(x)" increases.


If "x>1, f'(x)<0, f(x)" decreases.


The function "f(x)" increases on "(-\dfrac{\sqrt{3}}{3}, \dfrac{\sqrt{3}}{3})."

The function "f(x)" decreases on "(-\infin, -\dfrac{\sqrt{3}}{3})\cup(\dfrac{\sqrt{3}}{3}, \infin)."


The function "f(x)" has a local maximum at "x=1."

The function "f(x)" has a local minimum at "x=-1."



"f''(x)=(1-x^2)'=-2x"




"f''(x)=0=>-2x=0"

"x=0"

"f(0)=0-\dfrac{(0)^3}{3}=0"

The function "f(x)" is concave up on "(-\infin, 0)."

The function "f(x)" is concave down on "(0,-\infin)."


The point of inflection is "(0,0)."